6. Solve the differential equation by using the method of undermined coefficients: y - 2y + y = 2e - 3e-
Answer #1

y - Qy ty: 2 ex- 3 ex. dry - 2 dy + y 2 0 2 - 3 e? dx dx I let d =D D?y - 2 Dyt y = 284-3 e-? [ (02-20 + 1) y = 26² - 3 exolet y = AX + Bere y = Aex - Be-x Y = AX + B ex putting all these in maino Differential eqh - 0 A ex+ Be-x) - 2 (Aen-Bex) + AHere with this quess we are not able to find arunce o we need to guess another o Particular integral with the help of cof. -4Another guess - yo - A ? ex Be-x у: A (12. e1t exy 27 ) - Ве - 1 0 1 = A (?ext exx 2 x) (ex, I-е) +ве - 7 ут: Ах? е? + Aa e12А 2 А - ч . . 3D В - 3 Уе - ? е - 3 е-х ч 9. с. бр. у (1x C ) е 1 xex 3